# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 1.2.3 Capabilities of 6G
1.3 Overview of AI and communication for 6G
@@ 1.3.1 Drivers of Integrating AI into Wireless Communication
Future wireless network systems will evolve into integrated wireless infrastructure platforms deeply integrating “communication,” “perception,” “computing,” “intelligence,” and “storage.” These platforms will be capable of providing customized or personalized services on demand, surpassing the capabilities of traditional communication systems that rely on fixed architectures and predefined rules. However, such systems will need to leverage AI technologies. They will utilize big data and knowledge bases for inference and decision-making, employing models’ generalization abilities to adapt to various environments and scenarios and providing optimal resource allocation and management solutions. AI-enabled wireless communication networks will primarily enhance the following capabilities:
Big data processing capabilities: Mobile communication networks generate vast and diverse data every moment[13, 14, 15]. The International Data Corporation has predicted that by 2025, the global daily data generation will reach 491 exabytes (EB). This data typically includes terminal, wireless air interface, network, and service data in various formats, such as text, images, XML, HTML, graphics, and audio/video information, as shown in Table 2. Network operators can use AI technologies to train and make inference decisions on different types of mobile big data, enhancing network performance and optimizing across various dimensions and objectives.
The essential advantage of AI lies in its ability to extract information from unstructured mobile data, obtain abstract correlations, achieve effective organization of complex data, and reduce preprocessing work. This process does not require manual operations, thereby lowering labor costs.
Table 2: Direction of mobile communication network data and application
@@
Wireless air interface data Channel state information, intra/inter-cell interference, multipath delay Channel state information prediction, wireless channel modeling, power control, interference cancellation
@@
Network data Key performance indicator, extended detection and response, running log Dynamic load balancing and interference avoidance in wireless networks, network energy conservation, intelligent routing
@@
Breaking performance limitations of hierarchical modules: These models optimize source and channel independently. This is based on infinite data packet length assumptions and static channel conditions. In practical systems, coding is often finite, and channels are time-varying. Therefore, in dynamic and complex real-world application scenarios, even if each module operates optimally, the overall information transmission performance of the communication system may not be optimal. The advent of AI provides a new approach to improving the optimization design methods of traditional model-driven wireless communication systems, enhancing system performance and adaptability.
@@
AI can theoretically learn hidden structures and parameters to fit arbitrarily complex functions. This provides a more effective way to sense the complex and variable wireless environment and characterize the network state space. On the other hand, the designs of different communication system modules might have conflicting goals, and there might be performance constraints between modules. There are trade-offs between performance metrics, such as channel capacity and interference, transmission reliability, and system energy consumption. Optimizing each module individually often fails to achieve overall optimal performance. In such cases, AI can facilitate joint optimization design among modules. By leveraging advanced AI technologies like deep learning, different modules can learn and optimize together, thus achieving a more optimized overall system performance.
@@
International communication standard organizations, such as ITU, 3GPP, and IMT-2030, have initiated the standardization of AI applications in communication systems, and significant progress has been made. In the future, AI will further empower the infrastructure within networks, network management and operation systems, and service systems, fully unleashing the potential of integrating wireless communication and AI in the communication ecosystem and applications.
@@ 1.3.2 Vision of AI and 6G Integration
The advancement of AI technology has significantly empowered the development of wireless networks. The application of AI to enhance wireless networks is generally referred to as AI for Network (AI4NET). Its essence lies in using AI to improve wireless networks’ performance, efficiency, and user experience. Research in AI4NET primarily focuses on utilizing AI to optimize traditional algorithms, enhance network functionalities, and improve network operations and management. Figure 3 illustrates the degree of integration between 6G and AI from three perspectives: AI for Network, Network for AI, and AI as a Service. AI4NET’s application in 5G networks has significantly facilitated the intelligent development of mobile communication networks and vertical industries. Its role primarily manifests in operational and maintenance intelligence and network element intelligence [16].
@@
The former emphasizes utilizing AI to optimize traditional algorithms and automate and intellectualize tasks across network layers and operations. For instance, AI empowers wireless air interface transmission to achieve adaptive CSI feedback, beam management, and other enhancements to transmission performance. The latter focuses on proposing AI-driven intelligence for individual network elements and functional entities, breaking down proactive network responses into predictions of multiple performance metrics. This approach ultimately aims to optimize network resource allocation and reduce network latency. Examples include AI-based traffic forecasting, mobility management, and anomaly detection applications. However, in 5G networks, AI4NET mainly operates in an “add-on” mode, introducing additional latency and data processing bottlenecks. This approach somewhat hinders AI applications’ effectiveness and can raise concerns regarding data privacy and security issues.
@@
AI’s evolution and developmental needs in wireless networks point towards 6G networks no longer serving as communication networks. Instead, they will transform into integrated information networks that combine communication, perception, and computing into intelligent services. 6G networks will further deepen the integration of AI with wireless networks. While AI empowers 6G networks, 6G networks will also empower AI, achieving Network for AI (NET4AI), which provides comprehensive support for the deployment and application of AI in wireless environments [17]. The essence of NET4AI is to provide various support capabilities through the network for AI, enabling more efficient and real-time AI training/inference and enhancing data security and privacy protection. An illustration of AI-native 6G has been shown in [18], which presents a vision of a new programmable network device architecture for the NET4AI phase. The authors of [19] proposed a novel network architecture for providing native support for AI, called an AI-oriented network, which is represented as a network management framework with distributed AI computational capabilities and multi-party participation built in 6G networks. The authors of [20] combined two emerging key technologies, DT and generative AI, to provide a new way of thinking about the intelligence of 6G networks. The authors of [17] briefly described the two phases of AI4NET and NET4AI at the link and network levels and argued that NET4AI will be the backbone of 6G and future wireless networks.
@@
Moreover, 6G networks are expected to incorporate an AI plane from its inception and other newly added logical planes such as computing, data, and security. These will be interfaced in the architecture design to support functionalities like AI plane connectivity, AI plane data operations, AI plane computing capabilities, and AI plane security features. This expansion moves beyond traditional network scopes of connectivity services to encompass capabilities in computing power, data handling, algorithms, and more. The goal is to fully embed the intelligence provided by AI into the interconnection functionality of 6G networks. Additionally, through the data flow capabilities provided by 6G networks and various vertical tasks, constructing and training new AI models’ architecture NET4AI aims to achieve native AI (Native AI).
@@
Furthermore, to effectively support “inherent intelligence” and achieve ubiquitous “Native AI,” 6G networks will treat AI as a service for provision and processing. This concept gives rise to the idea of 6G AI as a Service (AIaaS). Specifically, 6G AIaaS utilizes resources and functionalities within the network (including 6G core networks, wireless access networks, and terminals), such as connectivity, computing, data, and models. It aims to construct a distributed, efficient, energy-efficient, and secure AI service ecosystem. This ecosystem includes AI model training, inference, deployment, and other functionalities in a low-carbon open environment. The transition from network-oriented AI in 5G to AI-oriented 6G networks aims to provide network experiences that surpass data-centric approaches and directly cater to user needs [21]. Not only can it redefine the ecosystem of edge cloud, but it can also build new business models through 6G mobile networks, transitioning from past connectivity-oriented networks to service-oriented networks, ultimately achieving ubiquitous intelligence [22, 23].
@@
AIaaS’s typical scenarios include but are not limited to, smart cities, smart agriculture, universal education, and smart industry. Typical applications include unmanned taxi services, smart grid inspections, home health monitoring, and virtual classrooms. The specific AI services involved can be categorized from different perspectives. On the one hand, from the perspective of object orientation, they can be divided into two main categories. One category provides AI services for 6G networks, which are used to optimize network performance, corresponding to AI4NET mentioned earlier. Examples include using end-to-end AI for customized air interface and network optimization and automated maintenance, leveraging AI trained on historical user data to predict traffic and optimize network resources. The other category provides AI services for end-users and third-party users (e.g., hospitals, autonomous vehicles, factories), corresponding to NET4AI, as previously discussed. For example, in industrial internet scenarios, wireless networks provide AIaaS to automatically identify industrial product quality using AI image-related technologies and trigger timely actions by industrial robots using reinforcement learning technologies. Each industrial detection scenario may require fine-tuning machine learning algorithms with a long deployment cycle. 6G networks enable rapidly deploying targeted AI services through large-scale distributed training, integration, and inference. On the other hand, AI services come in various types. According to the different stages of the AI model’s lifecycle, they can typically be classified into data-oriented AI services, training-oriented AI services, inference-oriented AI services, and validation-oriented AI services. These AI services vary in emphasis, quality assessment metrics, and key performance indicators.
@@
The Quality of AI Service (QoAIS) is a multidimensional evaluation and assurance framework within AIaaS, quantifying and grading user demands for AI service quality at the user level. QoAIS integrates into the intelligent native network of 6G to guide the provision of corresponding AIaaS and ensure user requirements are met. However, QoAIS cannot directly adopt traditional QoS metrics from communication networks, which primarily focus on performance aspects of connectivity services such as transmission rates and latency. With the evolution of 6G networks, alongside traditional communication resources, multidimensional resources will be integrated, including storage, distributed heterogeneous computing capabilities, data, and AI algorithms. Therefore, assessing the quality of native AI services within the network requires consideration across multiple dimensions, including connectivity, computing capabilities, algorithms, and data. As concerns over data privacy protection and network autonomy grow within the industry, the design of the QoAIS metric system should also incorporate dimensions such as security, privacy protection, and autonomous management [24].
@@
Moreover, with the continuous advancement of AI technology, Large Language Models (LLMs) represent a significant breakthrough in the field of NLP, contributing to the development of 6G AI. Compared to traditional small-parameter models, LLMs demonstrate robust contextual understanding, coherent text generation, logical reasoning, and generalization capabilities. Existing universal large models, such as GPT, can handle various domain-specific problems, demonstrating competence in addressing knowledge and issues within the communication domain as well [25]. The emergence of LLMs can bring a new paradigm to integrating 6G and AI. The intrinsic AI architecture in 6G can provide linkage, computation, model decomposition, and model distribution services for the training and inference of LLMs. LLMs can empower various domains of 6G networks (e.g., air interface, network side, network security, network operation and maintenance, etc.), enabling intelligent services and closed-loop control. In 6G networks, customized specialized large models can be deployed across different network layers. These specialized large models can collaboratively address network issues through task decomposition and composition, cross-layer collaboration, cloud-edge collaboration, and other modes, thereby enhancing the intelligence level of the network.
@@
2 History and Development of AI
2.1 Traditional AI Techniques
In 1950, Alan Turing introduced the historically influential Turing test in the paper [26], establishing a theoretical framework and foundational ideas of AI. This test is a benchmark for evaluating machine intelligence. The convening of the Dartmouth Conference in 1956 marked a watershed moment in the history of AI, and researchers reached a landmark consensus: all forms of learning behavior and intellectual characteristics can be thoroughly understood and analyzed in theory. These studies advanced AI in the 1950s and 1960s [27]. Firstly, game theory emerged as a cornerstone for studying intelligent strategies, highlighting AI’s prowess in strategic planning and handling complex scenarios. Computer scientists utilized advanced programming techniques to enable computers to compete against humans in strategy games, enhancing public awareness of AI’s potential and driving further algorithm design and decision theory developments. Secondly, early explorations into NLP made notable strides. Despite its infancy, initial machine translation systems began employing rule-based matching and simple statistical models to achieve automatic translation between different languages, showing AI’s vast potential in language understanding and semantic analysis. Additionally, AI has made remarkable achievements in logic and mathematics. Particularly in automated theorem proving, computers demonstrated formidable capabilities in automated reasoning and theorem verification through intricate symbolic logic systems and intelligent inference algorithms.
In the 1980s and 1990s, the emergence of expert systems[28] once again sparked a surge of research in AI. These systems emulated the decision-making processes of human experts in domains such as chemical analysis, medical diagnosis, and logical reasoning, thereby demonstrating the practical value and diverse applications of AI. These early expert systems achieved breakthroughs in their respective specialized domains and accumulated rich experience for subsequent AI research. Since 2006, deep learning technology [29] opened up new directions for the development of intelligent applications. In 2012, in the ImageNet Large Scale Visual Recognition Challenge, the AlexNet model[30] achieved remarkable performance and won the championship. AlexNet showed the enormous potential of CNNs in handling complex image data.
2.1.1 Machine Learning
Machine learning [31] empowers computers to learn and improve from data without explicit programming of detailed instructions, understanding the underlying patterns and regularities within data. As shown in Figure 5, machine learning can be primarily categorized into supervised, unsupervised, and reinforcement learning. Typical algorithms are summarized in Table 3. Supervised learning[32] relies on labeled data to guide the learning process. Unsupervised learning [33] learns the knowledge from datasets without labels. Reinforcement learning[34], as an advanced computational paradigm, draws inspiration from the learning processes of biological organisms in natural environments.
Figure 5: Relationship between AI and algorithms
Table 3: Typical algorithms of machine learning
Category Algorithm Description
Supervised Learning Linear Regression Predict the linear relationship
Decision Trees Tree based decision-making
Random Forest Integrate multiple decision trees
K-Nearest Neightbors Classify samples based on distance
Support Vector Machine Segmenting data through hyperplanes
Unsupervised Learning Cluster Analysis Grouping data based on similarity
Association Rule Learning Discover the dependencies in data
Auto-Encoder Learn the feature expression
Principal Component Analysis Extract the features of data
Reinforcement Learning Q-learning Optimize strategy by Q-table
Sarsa Optimize strategy by current strategy
Actor-Critic Combine gradient and value function
2.1.2 Deep Learning
The principle of deep learning technology is to emulate the human brain [35], constructing artificial neural networks to learn knowledge from data. The neural layers can automatically learn and extract complex feature patterns from large-scale datasets, using gradient descent to update parameter errors. Table 4 summarizes some popular deep learning algorithms, and detailed descriptions are given as follows. Convolutional neural network (CCN) are widely used for image and video understanding [36]. Inspired by the functioning principles of the biological visual system, CNNs efficiently extract local features from image data via convolution operations and reduce computational complexity through pooling layers. Recurrent Neural Networks (RNN) represent a class of neural networks designed explicitly for sequential data within deep learning. Leveraging a recursive network structure[37], RNNs excel at capturing latent dependencies in sequential data. Generative adversarial networks (GANs) are a class of neural networks widely used for unsupervised learning. They comprise two neural networks, a discriminator and a generator, and use adversarial training to produce artificial data identical to actual data [38]. Deep Reinforcement Learning (DRL) is a convergence of deep learning and reinforcement learning[34], employing deep neural networks to approximate value functions or policy functions in reinforcement learning problems. This approach enables agents to autonomously learn and optimize decision-making processes to maximize long-term rewards in complex and dynamic environments[39, 40].
Table 4: Typical algorithms of deep learning
Category Typical Algorithm Description
CNN LeNet-5 Early convolutional neural network
AlexNet Introduce ReLU and GPU acceleration
ResNet Introduce residual connections
RNN LSTM Introduce gate mechanism
GRU Simplified version of LSTM
Seq2Seq Translation and Abstract Model
GAN WGAN Introducing Wasserstein distance
CycleGAN Image to image without paired data
StyleGAN Generate high-quality images
DRL DQN Solve high-dimensional state problems
DDPG RL for continuous action spaces
PPO Improvement of Policy Gradient
2.2 Overview of Large Language Models
2.2.1 Generative Models
Figure 6: History of generative neural models
Generative models [41] learn the distribution of data models to generate new data. Figure 6 outlines the history of generative models. Hidden Markov models [42] and Gaussian mixture models [43] were proposed in the 1950s for continuous data like speech and sequences. However, the advent of deep learning significantly enhanced the performance of generative models. In NLP, traditional N-gram models[44] struggled with long sentences, leading to the adoption of RNN [45] for modeling longer dependencies. The development of LSTM [46] and GRU [47] further enhanced performance by controlling memory during training.
Generative adversarial networks (GANs) [48] were proposed in 2014. Besides GANs, other methods like Variational Autoencoders [49], and diffusion models have been developed to provide finer control over the image generation process and produce high-quality images. The development paths of generative models for NLP tasks and images differ across fields but eventually converge on the Transformer model architecture. Introduced by Vaswani et al. in 2017, the Transformer model revolutionized NLP tasks and was subsequently applied to computer vision, becoming the dominant architecture for many generative models[50],[51]. Prominent large language models such as BERT and GPT use the Transformer architecture as their primary building block, outperforming traditional models like LSTM and GRU. In computer vision, models like the Vision Transformer [52] and Swin Transformer[53] advanced this concept by integrating Transformer models with visual components for image-based tasks. This cross-domain convergence facilitated the development of multimodal models like CLIP [54], integrating transformer architecture with visual components for image encoding and multimodal prompt generation. Large language models such as GPT-3 achieved state-of-the-art performance on various natural language processing tasks. Many large models were also introduced, including Stable Diffusion[55] and the Sora.
2.2.2 Datasets of Large Language Models
Datasets play a pivotal role in the training of LLMs. The selection and construction of datasets are critical factors in ensuring that models can learn and generalize effectively. Some popular datasets are outlined as follows.
In early studies, ImageNet[56] is a prominent large-scale image database from Stanford University. It contains over 14 million images spanning diverse categories, including animals and plants. ImageNet serves as a benchmark dataset for evaluating the performance of image classification. COCO dataset[57] is built by Microsoft. It includes over 330,000 carefully curated images, with detailed annotations for 220,000 of these images, covering 80 different object categories. COCO is treated as an essential benchmark for evaluating image processing algorithms. BookCorpus [58] is a widely used multilingual book corpus that systematically collects and formats publicly available e-book resources, providing a large-scale text database that is accessible and easy to manipulate for both academia and industry. Pile dataset[59] is a comprehensive open-text resource library that integrates over 20 different types of subsets, including literary works, news articles, online encyclopedias, programming code, academic papers, and dialogue transcripts. Audio Set[60], meticulously constructed by Google’s research, is a vast audio recognition dataset that has become an essential resource for advancing sound recognition, classification, and event detection. HowTo100M[61] dataset aggregates one million instructional videos from YouTube, covering various topics, including cooking, crafts, and personal care.
2.2.3 Large Language Models
The development of large language models (LLMs) has been astonishingly rapid. Early models like T5[62] had only around a billion parameters, but with the application of scaling laws, recent models such as PaLM have grown to hundreds of billions of parameters. This highlights the rapid expansion in the scale of LLMs. Regarding performance, early-stage LLMs have already dominated conventional NLP tasks, outperforming other models. Notably, the recent GPT-4 model[63] can match or even surpass human performance in various exams, such as the US Bar Exam[64].
In 2020, OpenAI introduced GPT-3, trained on over 45TB of text data with an astounding 175 billion parameters. In early 2022, InstructGPT and GPT-3.5 were released, both large language models based on GPT-3. InstructGPT was designed to execute specific instructions, while GPT-3.5 aimed to align more closely with human values. Then, in November 2022, OpenAI launched ChatGPT, a conversational model that builds upon the GPT series. Besides the GPT series, other notable open-source large models include Llama-2-70B[65] and Zephyr-7B[66]. Llama-2-70B, trained on a vast dataset of 20 trillion tokens, excels in general benchmarks, with its variant Llama-2-chat-70B achieving a remarkable 92.66% win rate in AlpacaEval[67], surpassing GPT-3.5-turbo, though GPT-4 leads with 95.28%. Zephyr-7B, using distilled preference optimization, scores 90.6% on AlpacaEval and outperforms Llama-2-chat-70B on MT-Bench[68]. Additionally, WizardLM-70B[69] scores 7.71 on MT-Bench, slightly below GPT-3.5-turbo and GPT-4. Despite Zephyr-7B’s strong MT-Bench performance, it ranks poorly at 52.15% on the open-source LLM Leaderboard. In contrast, GodziLLa2-70B combines Maya Philippines datasets with Llama-2-70B, achieving a competitive 67.01%, akin to GPT-3.5-turbo but below GPT-4’s 85.36%. UltraLlama matches GPT-3.5-turbo’s performance and excels in specialized domains. While not matching proprietary counterparts, these open-source models provide robust tools and resources for advancing natural language processing, driven by community efforts and expanding AI technology applications.
A line of large vision models has been proposed for images. DALL-E 2[70] is a powerful image generation model developed by OpenAI, incorporating multiple advanced technologies, including the CLIP model’s text-to-image alignment capabilities, diffusion model’s generation mechanism, and the Transformer architecture’s deep learning capabilities. Google’s Parti[71] is a text-to-image generation model. It decomposes images into multiple image tokens and employs ViT-VQGAN technology for efficient encoding. This process involves a sizeable encoder-decoder network with 20 billion parameters. Google’s Muse[62] is a text-to-image generation Transformer model and utilizes a masked transformer architecture to perform efficient masking tasks in a discrete token space.
We also witnessed a series of video models. OpenAI’s SORA [72] is a revolutionary text-to-video generation tool that has opened new frontiers in the creative industry. This model rapidly transforms textual descriptions into visual stories, providing an unprecedented creative tool for professionals with agility, intelligence, and adaptability across various fields. Luma AI’s Dream Machine is an advanced video generation tool that employs a flexible and efficient multi-modal Transformer architecture trained directly on video data. This model can produce high-quality, realistic video content and simulate physical phenomena such as gravity, collisions, and light variations, ensuring visual and physical coherence. Google’s Veo is a large-scale video generation model that integrates several state-of-the-art technologies, including GQN, DVD-GAN, Imagen-Video, and Phenaki, and advanced models such as Transformer and Gemini. Veo can generate high-quality video content from various input modalities, including text, images, and videos.

### v3

@@ 1.2.3 Capabilities of 6G
Among the newly added capabilities mentioned above, the applicable AI-related capabilities draw the most attention. The possession of AI-related capabilities by 6G networks can be understood from several aspects. Firstly, it implies that the network can optimize its own performance through AI. Secondly, it indicates that the network is capable of providing support for the operation of AI. The ultimate form will be that the network can offer AI services to users and equipment just as it provides communication connections. Such capabilities can only be achieved through the in-depth fusion of wireless communication and AI.
1.3 Overview of AI and Communication Integration in 6G
In the vision and new capabilities of 6G network, the integration of AI and wireless communication is emphasized repeatedly. In this context, this section will elaborate on why and how to integrate AI and wireless communication. Generally speaking, AI will endow wireless networks with “autonomy” and “intelligence”, while wireless networks will provide AI with a broad application space, forming a two-way cyclic enabling. These two inseparable strategic development fields will continuously promote the convergent innovation and technological innovation of information, communication, and computing, and realize people’s visionary prospects in the spatiotemporal domain, information interaction types, and cross-industry convergent and innovative applications.
@@ 1.3.1 Drivers of Integrating AI into Wireless Communication
Future wireless network systems will evolve into integrated wireless infrastructure platforms deeply integrating “communication,” “perception,” “computing,” “intelligence,” and “storage.” These platforms will be capable of providing customized or personalized services on demand, surpassing the capabilities of conventional communication systems that rely on fixed architectures and predefined rules. However, such systems will need to leverage AI technologies. They will utilize big data and knowledge bases for inference and decision-making, employing models’ generalization abilities to adapt to various environments and scenarios and providing optimal resource allocation and management solutions. This section will analyze the driving forces of the integration of wireless communication and AI from the following two perspectives.
1) The diverse capabilities of AI are expected to propel the performance limits of wireless communication systems to new heights. AI possesses remarkable capabilities in handling big data. Mobile communication networks generate vast and diverse data every moment[21, 22]. The international data corporation has predicted that by 2025, the global daily data generation will reach 491 exabytes [23]. This data typically includes terminal, wireless air interface, network, and service data in various formats, such as text, images, extensible markup language (XML), hyperText markup language (HTML), graphics, and audio/video information, as shown in Table 1. Network operators can use AI technologies to train and make inference decisions on different types of mobile big data, enhancing network performance and optimizing across various dimensions and objectives.
Table 1: Data Information Table
@@
Air interface data Channel state information, intra/inter-cell interference, multipath delay Channel state information prediction, wireless channel modeling, power control, interference cancellation
@@
Network data KPI, extended detection and response, running log Dynamic load balancing and interference avoidance in wireless networks, network energy conservation, intelligent routing
@@
Besides, AI has predictive capabilities that are absent in conventional algorithms. AI can learn from historical data via ML algorithms and construct models to forecast future network requirements and possible issues. In wireless communication, this indicates the capacity to predict network congestion, device failures, or alterations in user behavior in advance. For instance, AI can anticipate potential traffic peaks during specific holidays or large-scale events and take resource allocation measures in advance, such as increasing the capacity of temporary BS or optimizing routing strategies, thereby effectively averting network breakdowns and enhancing user experience.
What’s more, the adaptive capacity is also a prominent characteristic of AI. The wireless communication environment is highly dynamic and is affected by various factors such as weather, terrain, buildings, and changes in the location of mobile devices. The adaptive ability of AI enables it to make rapid adjustments according to environmental changes and fluctuations in network conditions. For example, when a user moves from indoors to outdoors or is in a high-speed moving vehicle, the signal strength and interference situation will change. AI can sense these changes in real time and automatically adjust communication parameters to ensure the stability of communication quality, adapt to different communication scenarios, and improve the reliability and flexibility of the wireless communication system.
@@
From another perspective, in contrast to the hierarchical management of different levels in conventional communication systems, AI can theoretically learn hidden structures and parameters to fit arbitrarily complex functions. This provides a more effective way to sense the complex and variable wireless environment and characterize the network state space. On the other hand, the designs of different communication system modules might have conflicting goals, and there might be performance constraints between modules. There are trade-offs between performance metrics, such as channel capacity and interference, transmission reliability, and system energy consumption. Optimizing each module individually often fails to achieve overall optimal performance. In such cases, AI can facilitate joint optimization design among modules. By leveraging advanced AI technologies like DL, different modules can learn and optimize together, thus achieving a more optimized overall system performance.
2) AI presents a novel and promising avenue for resolving the aforementioned challenges encountered in wireless communications. The core reason for the challenges summarized in Section 1.1 of this paper is that the current wireless network are not intelligent enough. Future wireless network systems will evolve into an integrated wireless infrastructure platform that deeply integrates “communication”, “sensing”, “computation”, “intelligence”, and “storage”. It will provide personalized services, which are far beyond the processing and management capabilities of conventional artificial rule pre-definition and execution. Therefore, communication systems require an automated and intelligent architecture and process to ensure the operation and development of networks and services. AI technology, which makes inferences and judgments based on big data and knowledge bases and utilizes the generalization ability of models to adapt to various environments and scenarios, can provide optimal resource allocation and management solutions and is the most promising way to achieve the above goals.
Specifically, on the wireless radio access network (RAN) side, AI can be combined to conduct data analysis and decision-making for the control unit plane of 5G new radio (NR), realizing the immediate dynamic performance optimization of wireless resource scheduling and management. The next generation node B (gNB) BS can achieve intelligent on/off switching based on AI to achieve dynamic energy savings. On the CN side, with the network data analysis function (NWDAF) network elements achieving comprehensive interconnection with surrounding CN elements and realizing data collection, AI will also participate in the control and decision-making of the CN, including quality of service (QoS) guarantee, traffic handing, 5G edge computing, and load balancing of network functions.
@@
International communication standard organizations, such as ITU, 3rd generation partnership project (3GPP), and IMT-2030, have initiated the standardization of AI applications in communication systems, and significant progress has been made. In the future, AI will further empower the infrastructure within networks, network management and operation systems, and service systems, fully unleashing the potential of integrating wireless communication and AI in the communication ecosystem and applications.
@@ 1.3.2 Vision of AI and 6G Integration
The advancement of AI technology has significantly empowered the development of wireless networks. The application of AI to enhance wireless networks is generally referred to as AI for Network (AI4NET). Its essence lies in using AI to improve wireless networks’ performance, efficiency, and user experience. Research in AI4NET primarily focuses on utilizing AI to optimize conventional algorithms, enhance network functionalities, and improve network operations and management. Figure 3 illustrates the degree of integration between 6G and AI from three perspectives: AI4NET, network for AI (NET4AI), and AI as a service (AIaaS). AI4NET’s application in 5G networks has significantly facilitated the intelligent development of mobile communication networks and vertical industries. Its role primarily manifests in operational and maintenance intelligence and network element intelligence [24].
@@
The former emphasizes utilizing AI to optimize conventional algorithms and automate and intellectualize tasks across network layers and operations. For instance, AI empowers wireless air interface transmission to achieve adaptive channel state information (CSI) feedback, beam management, and other enhancements to transmission performance. The latter focuses on proposing AI-driven intelligence for individual network elements and functional entities, breaking down proactive network responses into predictions of multiple performance metrics. This approach ultimately aims to optimize network resource allocation and reduce network latency. Examples include AI-based traffic forecasting, mobility management, and anomaly detection applications. However, in 5G networks, AI4NET mainly operates in an “add-on” mode, introducing additional latency and data processing bottlenecks. This approach somewhat hinders AI applications’ effectiveness and can raise concerns regarding data privacy and security issues.
@@
AI’s evolution and developmental needs in wireless networks point towards 6G networks no longer serving as communication networks. Instead, they will transform into integrated information networks that combine communication, perception, and computing into intelligent services. 6G networks will further deepen the integration of AI with wireless networks. While AI empowers 6G networks, 6G networks will also empower AI, achieving NET4AI, which provides comprehensive support for the deployment and application of AI in wireless environments [25]. The essence of NET4AI is to provide various support capabilities through the NET4AI, enabling more efficient and real-time AI training/inference and enhancing data security and privacy protection. An illustration of AI-native 6G has been shown in [26], which presents a vision of a new programmable network device architecture for the NET4AI phase. The authors of [27] proposed a novel network architecture for providing native support for AI, called an AI-oriented network, which is represented as a network management framework with distributed AI computational capabilities and multi-party participation built in 6G networks. The authors of [28] combined two emerging key technologies, DT and generative AI, to provide a new way of thinking about the intelligence of 6G networks. The authors of [25] briefly described the two phases of AI4NET and NET4AI at the link and network levels and argued that NET4AI will be the backbone of 6G and future wireless networks.
@@
Moreover, 6G networks are expected to incorporate an AI plane from its inception and other newly added logical planes such as computing, data, and security. These will be interfaced in the architecture design to support functionalities like AI plane connectivity, AI plane data operations, AI plane computing capabilities, and AI plane security features. This expansion moves beyond conventional network scopes of connectivity services to encompass capabilities in computing power, data handling, algorithms, and more. The goal is to fully embed the intelligence provided by AI into the interconnection functionality of 6G networks. Additionally, through the data flow capabilities provided by 6G networks and various vertical tasks, constructing and training new AI models’ architecture NET4AI aims to achieve native AI.
@@
Furthermore, to effectively support “inherent intelligence” and achieve ubiquitous “Native AI,” 6G networks will treat AIaaS for provision and processing. This concept gives rise to the idea of 6G AIaaS. Specifically, 6G AIaaS utilizes resources and functionalities within the network (including 6G CNs, wireless access networks, and terminals), such as connectivity, computing, data, and models. It aims to construct a distributed, efficient, energy-efficient, and secure AI service ecosystem. This ecosystem includes AI model training, inference, deployment, and other functionalities in a low-carbon open environment. The transition from network-oriented AI in 5G to AI-oriented 6G networks aims to provide network experiences that surpass data-centric approaches and directly cater to user needs [29]. Not only can it redefine the ecosystem of edge cloud, but it can also build new business models through 6G mobile networks, transitioning from past connectivity-oriented networks to service-oriented networks, ultimately achieving ubiquitous intelligence [30, 31].
@@
AIaaS’s typical scenarios include but are not limited to, smart cities, smart agriculture, universal education, and smart industry. Typical applications include unmanned taxi services, smart grid inspections, home health monitoring, and virtual classrooms. The specific AI services involved can be categorized from different perspectives. On the one hand, from the perspective of object orientation, they can be divided into two main categories. One category provides AI services for 6G networks, which are used to optimize network performance, corresponding to AI4NET mentioned earlier. Examples include using end-to-end AI for customized air interface and network optimization and automated maintenance, leveraging AI trained on historical user data to predict traffic and optimize network resources. The other category provides AI services for end-users and third-party users (e.g., hospitals, autonomous vehicles, factories), corresponding to NET4AI, as previously discussed. For example, in industrial internet scenarios, wireless networks provide AIaaS to automatically identify industrial product quality using AI image-related technologies and trigger timely actions by industrial robots using reinforcement learning (RL) [32] technologies. Each industrial detection scenario may require fine-tuning ML algorithms with a long deployment cycle. 6G networks enable rapidly deploying targeted AI services through large-scale distributed training, integration, and inference. On the other hand, AI services come in various types. According to the different stages of the AI model’s lifecycle, they can typically be classified into data-oriented AI services, training-oriented AI services, inference-oriented AI services, and validation-oriented AI services. These AI services vary in emphasis, quality assessment metrics, and key performance indicators (KPI).
@@
The quality of AI service (QoAIS) is a multidimensional evaluation and assurance framework within AIaaS, quantifying and grading user demands for AI service quality at the user level. QoAIS integrates into the intelligent native network of 6G to guide the provision of corresponding AIaaS and ensure user requirements are met. However, QoAIS cannot directly adopt conventional QoS metrics from communication networks, which primarily focus on performance aspects of connectivity services such as transmission rates and latency. With the evolution of 6G networks, alongside conventional communication resources, multidimensional resources will be integrated, including storage, distributed heterogeneous computing capabilities, data, and AI algorithms. Therefore, assessing the quality of native AI services within the network requires consideration across multiple dimensions, including connectivity, computing capabilities, algorithms, and data. As concerns over data privacy protection and network autonomy grow within the industry, the design of the QoAIS metric system should also incorporate dimensions such as security, privacy protection, and autonomous management [33].
@@
Moreover, with the continuous advancement of AI technology, large language models (LLM) [34] represent a significant breakthrough in the field of NLP, contributing to the development of 6G AI. Compared to conventional small-parameter models, LLM demonstrates robust contextual understanding, coherent text generation, logical reasoning, and generalization capabilities. Existing universal large models, such as GPT, can handle various domain-specific problems, demonstrating competence in addressing knowledge and issues within the communication domain as well [35]. The emergence of LLM can bring a new paradigm to integrating 6G and AI. The intrinsic AI architecture in 6G can provide linkage, computation, model decomposition, and model distribution services for the training and inference of LLM. LLM can empower various domains of 6G networks (e.g., air interface, network side, network security, network O&M, etc.), enabling intelligent services and closed-loop control. In 6G networks, customized specialized large models can be deployed across different network layers. These specialized large models can collaboratively address network issues through task decomposition and composition, cross-layer collaboration, cloud-edge collaboration, and other modes, thereby enhancing the intelligence level of the network.